# Learning atomic forces from uncertainty-calibrated adversarial attacks

### v2

@@ I Main
While such adversarial approaches have demonstrated great potential in active learning, very little attention has so far been put towards quantifying the errors of MLIPs on the adversarial structures. As such, a fundamental question remains: To what extent can we control the actual errors of the MLIP on the adversarial structures? To answer this question, we have developed the Calibrated Adversarial Geometry Optimization (CAGO) algorithm, which aims to discover new adversarial structures with target force errors preassigned by the user (Figure 1). To unify estimated prediction uncertainties with real errors, we perform uncertainty calibration. To control the MLIP prediction errors on adversarial structures, we optimize structures to moderate target errors. These structures are within the range of validity of the uncertainty calibration while being challenging structures for the MLIPs, from which they can learn. To demonstrate the usefulness of our approach, we integrate CAGO into an active learning framework, as shown in Figure 1, enabling us to learn liquid-water dynamics and water adsorption in metal-organic frameworks from small datasets.
@@ II.2 Adversarial structures
Like in ref. 23, we create adversarial attacks by optimizing the structure $\mathbf{x}$ according to a fitness function $\mathcal{L}$. However, instead of maximizing the committee uncertainty estimate $\hat{\sigma}$, we optimize calibrated prediction uncertainties $\sigma_{\text{cal}}$ towards the uncertainty target $\delta$:
@@ II.5 Calibrated adversarial geometry optimization to target error
Next, we perform CAGO to solve equation (8) without a bias-term by geometry optimization using numerical finite-difference gradients for 40 structures (see Figure 1, all structures reaching the target force uncertainty indicated in green). Figure 2c reports the corresponding error statistics, i.e., individual model forces versus reference DNN@MB-pol forces. The error statistics for adversarial structures for CAGO with calibrated uncertainties lead to a mean error close to the target error, whereas the uncalibrated case has errors about twice as high as the target. The potential usefulness of such calibrated adversarial structures can be understood when compared to ordinary MD structures, which have about half the error and would, therefore, not provide the same level of learning content if added to the training set.
Figure 3: Benchmark on the effect of CAGO algorithm hyperparameters. The panels show the distribution of MLIP force errors on adversarial structures, with corresponding mean errors in the inset of each panel. Line legends and colors are specified by ticks in inset figures. a. Force error statistics for different uncertainty targets. b. Force error statistics for adversarial structures optimized for a target 100 meV/Å uncertainty with different committee sizes, using uncertainty calibration from a committee of 20 MLPs. c. Force error statistics for adversarial structures optimized for a target 100 meV/Å uncertainty, considering uncertainty calibration performed with its respective committee size.
@@ II.6 Learning liquid water from a single structure
An MLIP is only as good as its reference data. Therefore, creating a training set with as high-quality references as possible is desirable. However, higher quality is generally associated with higher computational cost. In particular, density-corrected DFT (DC-DFT) and density-corrected R2SCAN (DC-R2SCAN) achieve excellent correspondence in energies with respect to coupled cluster methods.[30, 31] However, DC-DFT and DC-R2SCAN are rather expensive compared to ordinary functionals, making the standard route of starting from a training set of ab initio MD trajectories prohibitively expensive. Taking this to the extreme, here we use CAGO-based active learning to learn the MLIP for DC-R2SCAN starting from only a single structure of liquid water (Figure 4a). Figure 4 reports our benchmark on the performance of MLIP iterations during active learning for Allegro with error target 100 meV/Å, and DeePMD with target 100 meV/Å and 200 meV/ÅMLIPs (refer to Methods IV.2 for additional details).
@@
Figure 4: Benchmark of convergence of liquid-water properties for CAGO-based active learning starting out from a single structure. We report data for the Allegro MLIP with target error 100 meV and DeePMD with two different target errors, 100 meV and 200 meV (see Methods IV.2 for more details). Each point in the graphs has been computed from 12 models using the subset of stable MLIPs, with three replica simulations each, for a total of 540 and 1800 Allegro and DeePMD 2 ns MD simulations, respectively. The error bars are the standard deviation between the different committee members, and the horizontal lines correspond to the final value of the property for the model. a. The liquid water box structure that was used to start the active learning training. Oxygen is represented in red, with hydrogen in white. b. Percentage stable MLIPs in accordance with the stability criteria in Methods IV.5.2. c. Average force error of MLIPs on the liquid water structures. d. Mean liquid water density. e. Self-diffusion coefficient. f. First coordination number (integral of the first radial distribution function peak).
@@
Figure 4b reports the percentage of stable models. At around 80 training-set structures, Allegro models all become stable, while the DeePMD models achieve stability at around 220 or 350 for the 200 meV/Å and 100 meV/Å targets, respectively. The slower convergence of DeePMD is in line with past research, which shows that MLIPs based on equivariant layers are not only more accurate but also more data-efficient,[2, 3] with similar equivariant NequIP also being benchmarked to be more stable than DeePMD.[6] However, DeePMD, with our CAGO structures, performs rather well compared to its data-hungry reputation.[2] To put this into context, ref. 6 used 10 000 structures and achieved stability for an average of 0.247 nanoseconds under $NVT$ conditions with a simple water force-field. In contrast, we achieve 100 % stability for two nanoseconds under the more challenging $NPT$ conditions for a complex many-body ab initio method.
@@
To benchmark how well CAGO-based active learning performs for energetics, we report in Figure 4c the average force error for 100 liquid water MD structures. In accordance with the literature,[3] the force errors for the trained models converge to different levels of error for Allegro (37 meV/Å) and DeePMD (63 meV/Å), with the Allegro force errors being smaller than those of DeePMD. Similar force error magnitudes were also reported in ref. 2. It is noteworthy that while CAGO distorts the structures from MD, force errors in undistorted liquid MD structures improve. In particular, for both targets with DeePMD, we see a similar error level. Given that 63 meV/Å is relatively close to the target 100 meV/Å, it may be more advantageous for DeePMD to use a higher target error than for Allegro, and we see that energetics do not suffer from this. In fact, the stability test shown in Figure 4b indicates that the 200 meV/Å target improves the stability convergence.
@@
Next, in Figures 4d, e, and f, we benchmark the thermodynamical, dynamical, and structural properties of liquid water, respectively. As with the stability, for all properties, Allegro converges around 80 structures, with DeePMD following closely. Overall, DeePMD exhibits larger variability for these properties. This is consistent with past benchmark studies on MB-pol-based DeePMD, which, even for huge training sets, struggle to achieve the correct density,[5, 32] whereas Allegro gets it spot on.[33] Nonetheless, we only observe variability to within a few percent.
@@ III Discussion
In this paper, we presented the CAGO algorithm to systematically generate adversarial molecular structures with user-assigned MLIP prediction errors. Through our benchmarks of the different hyperparameters of CAGO, we establish suitable minimums. In particular, we found it important to use many models, about 10 models in our case, to obtain a good error calibration, while the CAGO algorithm itself works well even for small committees. For the true error to converge to the target error, moderate uncertainty targets should be used. This is in line with our error calibration analysis, where we saw that error calibration is less robust for the high-error regime, where calibration data is scarce.
These insights enabled us to perform active learning using minimal datasets with simple active learning protocols. In particular, we learned liquid water from a single structure of water for both Allegro and DeePMD and only 11 for water adsorption in the UiO-66 metal-organic framework. Although CAGO seemingly introduces an extra step of complexity into the active learning scheme, it ultimately simplifies the rest of the workflow. With CAGO, fewer manual adjustments and tricks are needed, such as low/high-temperature sampling or reducing the size of MD steps to achieve a stable model. Moreover, we have used orders of magnitude fewer structures than is typically reported in other works on adversarial active learning.[23, 24] A key problem with current MLIPs is their typical dependence on large and approximate but cheap DFT datasets. CAGO-based active learning offers a promising solution by only requiring a few structures, allowing highly accurate but costly coupled-cluster and quantum Monte Carlo reference calculations to be used as training data.
@@ IV.5.1 Calibrated adversarial geometry optimization to target error
For benchmarking the CAGO algorithm in section II.5, we targeted the DNN@MB-pol water model as specified in Methods IV.3. First, we trained a committee of 20 DeePMD MLIPs on a training set of MD trajectories. Second, using these 20 models and a separate test set, we performed uncertainty calibration in accordance with equation (5), the results of which are reported in Figure 2. Third, picking out 40 structures from the test set, we perform CAGO with various hyperparameters, such as different error targets and committee sizes. The errors we report are root mean square errors of individual models against DNN@MB-pol. For more details on the training and test set, we refer to the Supplementary Information.
@@ IV.5.2 Learning liquid water from a single structure
For the active learning workflow, we used committees with 12 MLIPs to learn liquid water from a single structure of 64 water molecules, with details on initial training sets reported in Supplementary Information. In each of the first two active learning iterations, we sampled five structures from the training set and optimized the structures with CAGO using random subsets of 3 committee members. For all subsequent iterations, MD sampling with LAMMPS [42] was performed before CAGO (see Section IV.5), and in the case of pure water, 20 samples were extracted and taken for CAGO considering all 12 committee members. Due to the low number of structures in the first iterations of active learning, the MLIPs tend to be unstable; we, therefore, performed $NVT$ simulations at 300 K to generate new structures for CAGO for iterations 3 and 4. These simulations were run for 50 ps using the velocity-Verlet integrator with a 0.25 fs timestep and Nosé-Hoover thermostat chain with 3 thermostats and a 0.5 ps relaxation time. From iteration 5 and onwards, we performed $NPT$ simulations at 300 K and 1 bar, sampling for 100 ps, using the same settings for the thermostat part, and a Nosé-Hoover barostat chain of 3 barostats with 1 ps relaxation time. Structures derived from CAGO with maximum forces (mean by committee members) higher than 40 eV/Å or maximum force from reference calculations higher than 30 eV/Å, or minimum distance lower than 0.75 Å, or with convergence warnings from CP2K, were filtered out. Note that the DeePMD MLIPs used slightly varying settings for the training throughout the active learning loop to avoid overfitting. In the first two iterations, we trained the models for 50 000 steps to avoid overfitting with small training sets. As more structures were added, we increased the number of training steps to 150 000 for the next two active learning iterations, and from iteration 5 and onwards, we used 400 000 training steps.
@@ VII Acknowledgments
The work was supported by the Research Council of Norway through the Centre of Excellence Hylleraas Centre for Quantum Molecular Sciences (Grant 262695) and the Young Researcher Talent grants 344993 and 354100, as well as by the EuroHPC Joint Undertaking (Grants EHPC-REG-2023R02-088, EHPC-REG-2023R03-146). Support was also received from the Centre for Advanced Study in Oslo, Norway, which funded and hosted the SLB Young CAS Fellow research project during the academic year of 23/24 and 24/25. Part of the simulations were performed on resources provided by Sigma2 — the Norwegian National Infrastructure for High-Performance Computing and Data Storage (grant numbers NN4654K and NS4654K).

### v3

@@ I Introduction
While such adversarial approaches have demonstrated great potential in active learning, very little attention has so far been put towards quantifying the errors of MLIPs on the adversarial structures. As such, a fundamental question remains: To what extent can we control the actual errors of the MLIP on the adversarial structures? To answer this question, we have developed the Calibrated Adversarial Geometry Optimization (CAGO) algorithm, which aims to discover new adversarial structures with target force errors preassigned by the user (Figure 1). To unify estimated prediction uncertainties with real errors, we perform uncertainty calibration. To control the MLIP prediction errors on adversarial structures, we optimize structures to reach moderate target errors. These structures are within the range of validity of the uncertainty calibration while being challenging structures for the MLIPs, from which they can learn. To demonstrate the usefulness of our approach, we integrate CAGO into an active learning framework, as shown in Figure 1, enabling us to learn liquid-water dynamics and water adsorption in metal-organic frameworks from small datasets.
@@ II.2 Adversarial structures
Like in ref. 23, we create adversarial attacks by optimizing the structure $\mathbf{x}$ according to a fitness function $\mathcal{L}$. However, instead of maximizing the committee uncertainty estimate $\hat{\sigma}$, we optimize calibrated prediction uncertainties $\sigma_{\text{cal}}$ towards the error target $\delta$:
@@ II.5 Calibrated adversarial geometry optimization to target error
Next, we perform CAGO to solve equation (8) without a bias-term by geometry optimization using numerical finite-difference gradients for 40 structures (see Figure 1, all structures reaching the target force error indicated in green). Figure 2c reports the corresponding error statistics, i.e., individual model forces versus reference DNN@MB-pol forces. The error statistics for adversarial structures for CAGO with calibrated uncertainties lead to a mean error close to the target error, whereas the uncalibrated case has errors about twice as high as the target. The potential usefulness of such calibrated adversarial structures can be understood when compared to ordinary MD structures, which have about half the error and would, therefore, not provide the same level of learning content if added to the training set.
Figure 3: Benchmark on the effect of CAGO algorithm hyperparameters. The panels show the distribution of MLIP force errors on adversarial structures, with corresponding mean errors in the inset of each panel. Line legends and colors are specified by ticks in inset figures. a. Force error statistics for different error targets. b. Force error statistics for adversarial structures optimized for a target 100 meV/Å uncertainty with different committee sizes, using uncertainty calibration from a committee of 20 MLPs. c. Force error statistics for adversarial structures optimized for a target 100 meV/Å uncertainty, considering uncertainty calibration performed with its respective committee size.
@@ II.6 Learning liquid water from a single structure
An MLIP is only as good as its reference data. Therefore, creating a training set with as high-quality references as possible is desirable. However, higher quality is generally associated with higher computational cost. In particular, density-corrected DFT (DC-DFT) and density-corrected R2SCAN (DC-R2SCAN) achieve excellent correspondence in energies with respect to coupled cluster methods.[30, 31] However, DC-DFT and DC-R2SCAN are rather expensive compared to ordinary functionals, making the standard route of starting from a training set of ab initio MD trajectories prohibitively expensive. It is therefore important to have a method that can start the active learning process with as few training structures as possible. Taking this to the extreme, here we use active learning to learn an MLIP for DC-R2SCAN starting from only a single structure of liquid water (Figure 4a). Figure 4 reports our benchmark on the performance of MLIP iterations during CAGO-based active learning for Allegro with error target 100 meV/Å (Allegro CAGO 100 meV/Å), and, for DeePMD, with targets 100 meV/Å and 200 meV/Å (CAGO 100 meV/Å and 200 meV/Å), standard active learning selecting maximum uncertainty structures from MD (max. uncertainty), and sampling random structures from MD at 500 K (random 500 K), please refer to Methods IV.5.2 for additional details. Note that for the first two iterations of CAGO-based active learning, we do CAGO from the first starting geometry, and not from MD sampled geometries (see Methods IV.5.2 for additional details).
@@
Figure 4: Benchmark of convergence of liquid-water properties for CAGO-based and standard active learning starting out from a single structure. We report data for the Allegro MLIP with target error 100 meV/Å and for DeePMD with target error 100 meV/Å, 200 meV/Å, as well as picking maximum uncertainty structure from MD and random structure sampling from 500 K MD simulations (see Methods IV.5.2 for more details). Each point in every graph has been computed from 12 models using the subset of stable MLIPs, with three replica MD simulations lasting 2 ns each, for a total of 540 Allegro and 3744 DeepMD simulations. The error bars are the standard deviation between the different committee members, and the horizontal lines correspond to the average final value of the property for Allegro CAGO. a. The liquid water box structure that was used to start the active learning training. Oxygen is represented in red, with hydrogen in white. b. Percentage of stable MLIPs in accordance with the stability criteria in Methods IV.5.2. c. Average force error of MLIPs on the liquid water structures. d. Mean liquid water density. e. Self-diffusion coefficient. f. First coordination number (integral of the first radial distribution function peak).
@@
Figure 4b reports the percentage of stable models. At around 80 training-set structures, all Allegro models become stable, while the CAGO-based DeePMD models achieve stability at around 220 or 350 for the 200 meV/Å and 100 meV/Å targets, respectively. The slower convergence of DeePMD is in line with past research, which shows that MLIPs based on equivariant layers are not only more accurate but also more data-efficient,[2, 3] with similar equivariant NequIP also being benchmarked to be more stable than DeePMD.[6] However, DeePMD, with our CAGO structures, performs rather well compared to its data-hungry reputation.[2] To put this into context, ref. 6 used 10 000 structures and achieved stability for an average of 0.247 nanoseconds under $NVT$ conditions with a simple water force-field. In contrast, we achieve 100 % stability for two nanoseconds under the more challenging $NPT$ conditions for a complex many-body ab initio method. Interestingly, active learning using more standard approaches (max. uncertainty and random 500 K) achieves stability quicker than CAGO-based active learning. This could be due to the sampling yielding higher uncertainty structures than CAGO, which targets moderate force errors. This would be in line with the previously observed trend of higher target errors leading to more stable models.
@@
To benchmark how well CAGO-based active learning performs for energetics, we report in Figure 4c the average force error for 100 undistorted liquid water MD structures. In agreement with the literature,[3] the Allegro force errors are considerably smaller than those of DeePMD, with Allegro CAGO 100 meV/Å converging to 37 meV/Å, and DeePMD CAGO 100 meV/Å and 200 meV/Å converging to 63 meV/Å and 70 meV/Å, respectively. Similar force error magnitudes were also reported in ref. 2. It is noteworthy that while CAGO distorts the structures from MD during training, CAGO improves force errors of MLIPs on undistorted liquid MD structures. Although stability for DeePMD is achieved with fewer training structures for standard active learning approaches, the force errors are higher: max. uncertainty gives an error of 85 meV/Å, (approximately 35 % higher than CAGO 100 meV/Å), and random 500 K gives an error of 77 meV/Å (approximately 22 % higher than CAGO 100 meV/Å).
@@
Next, in Figures 4d, e, and f, we benchmark the thermodynamical, dynamical, and structural properties of liquid water, respectively. As with the stability, for all properties, Allegro converges around 80 structures, with DeePMD following closely. Overall, DeePMD exhibits larger variability for these properties. This is consistent with past benchmark studies on MB-pol-based DeePMD, which, even for huge training sets, struggle to achieve the correct density,[5, 32] whereas Allegro gets it spot on.[33] Nonetheless, we only observe variability within a few percent, and CAGO-based DeepMD models are within one standard deviation of the final Allegro iteration. For DeePMD, comparing against more standard active learning approaches (max. uncertainty and random 500 K), CAGO with 100 meV/Å performs best, followed by CAGO with 200 meV/Å, followed closely by random 500 K, and max. uncertainty. This order of accuracy is the same as the order of accuracy for the energetics as reported in Figure 4c (see Table S3 in the Supplementary Information for the numbers of the final models).
@@ III Discussion
In this paper, we presented the CAGO algorithm to systematically generate adversarial molecular structures with user-assigned MLIP prediction errors. Through our benchmarks of the different hyperparameters of CAGO, we establish suitable minimums. In particular, we found it important to use many models, about 10 models in our case, to obtain a good error calibration, while the CAGO algorithm itself works well even for small committees. For the true error to converge to the target error, moderate error targets should be used. This is in line with our error calibration analysis, where we saw that error calibration is less robust for the high-error regime, where calibration data is scarce.
These insights enabled us to perform active learning using minimal datasets with simple active learning protocols. In particular, we learned liquid water from a single structure of water for both Allegro and DeePMD and only 11 for water adsorption in the UiO-66 metal-organic framework. Although CAGO seemingly introduces an extra step of complexity into the active learning scheme, it ultimately simplifies the rest of the workflow. With CAGO, fewer manual adjustments and tricks are needed, such as low/high-temperature sampling or reducing the size of MD steps to achieve accurate models. Moreover, we have used orders of magnitude fewer structures than is typically reported in other works on adversarial active learning.[23, 24] A key problem with current MLIPs is their typical dependence on large and approximate but cheap DFT datasets. CAGO-based active learning offers a promising solution by only requiring a few structures, allowing highly accurate but costly coupled-cluster and quantum Monte Carlo reference calculations to be used as training data.
@@ IV.5.1 Calibrated adversarial geometry optimization to target error
For benchmarking the CAGO algorithm in section II.5, we targeted the DNN@MB-pol water model as specified in Methods IV.3. First, we trained a committee of 20 DeePMD MLIPs on a training set of MD trajectories. Second, using these 20 models and a separate test set, we performed uncertainty calibration in accordance with equation (5), the results of which are reported in Figure 2. Third, sampling 40 structures from the test set, we perform CAGO with various hyperparameters, such as different error targets and committee sizes. The errors we report are root mean square errors of individual models against DNN@MB-pol. For more details on the training and test set, we refer to the Supplementary Information.
@@ IV.5.2 Learning liquid water from a single structure
For the active learning workflow, we used committees with 12 MLIPs to learn liquid water from a single structure of 64 water molecules, with details on initial training sets reported in Supplementary Information. In each of the first two active learning iterations with CAGO, we sampled five structures from the training set and optimized the structures with CAGO using random subsets of 3 committee members. For all subsequent iterations, MD sampling with LAMMPS [42] was performed before CAGO (see Section IV.5), and in the case of pure water, 20 samples were extracted and taken for CAGO considering all 12 committee members. Due to the low number of structures in the first iterations of active learning, the MLIPs tend to be unstable; we, therefore, performed $NVT$ simulations at 300 K to generate new structures for CAGO for iterations 3 and 4. These simulations were run for 50 ps using the velocity-Verlet integrator with a 0.25 fs timestep and Nosé-Hoover thermostat chain with 3 thermostats and a 0.5 ps relaxation time. From iteration 5 and onwards, we performed $NPT$ simulations at 300 K and 1 bar, sampling for 100 ps, using the same settings for the thermostat part, and a Nosé-Hoover barostat chain of 3 barostats with 1 ps relaxation time. Structures derived from CAGO with maximum forces (mean by committee members) higher than 40 eV/Å or maximum force from reference calculations higher than 30 eV/Å, or minimum distance lower than 0.75 Å, or with convergence warnings from CP2K, were filtered out. Note that the DeePMD MLIPs used slightly varying settings for the training throughout the active learning loop to avoid overfitting. In the first two iterations, we trained the models for 50 000 steps to avoid overfitting with small training sets. As more structures were added, we increased the number of training steps to 150 000 for the next two active learning iterations, and from iteration 5 and onwards, we used 400 000 training steps. For active learning with maximum uncertainty and elevated temperature, MD-sampled structures were used for all iterations. 500 K was used for elevated temperature active learning. For maximum uncertainty, we picked structures based on the highest force uncertainty on any atom in the structure. For elevated temperature active learning, we sampled structures at random. In both cases we filter high-force structures as in the CAGO-based active learning.
@@ VII Acknowledgments
We would like to thank the anonymous reviewer for his/her suggestion of adding a comparison with standard active learning based on picking maximum uncertainty structures from MD, and random structures from elevated temperature MD. This work was supported by the Research Council of Norway through the Centre of Excellence Hylleraas Centre for Quantum Molecular Sciences (Grant 262695) and the Young Researcher Talent grants 344993 and 354100. We acknowledge the EuroHPC Joint Undertaking for awarding this project access to the EuroHPC supercomputer LUMI, hosted by CSC (Finland) and the LUMI consortium through a EuroHPC Regular Access call (Grants EHPC-REG-2023R02-088, EHPC-REG-2023R03-146). Support was also received from the Centre for Advanced Study in Oslo, Norway, which funded and hosted the SLB Young CAS Fellow research project during the academic year of 23/24 and 24/25. Part of the simulations were performed on resources provided by Sigma2 — the Norwegian National Infrastructure for High-Performance Computing and Data Storage (grant numbers NN4654K and NS4654K).